# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ Abstract.
Key words and phrases:
agreement digraph, agreement forest, phylogenetic network, SNPR distance, tree-child
All authors thank the New Zealand Marsden Fund for their financial support. Part of this paper is based upon work supported by the National Science Foundation under Grant No. DMS-1929284 while the second and third authors were in residence at the Institute for Computational and Experimental Research in Mathematics in Providence, Rhode Island, US, during the Theory, Methods, and Applications of Quantitative Phylogenomics program.
1. Introduction
@@
Phylogenetic trees and networks play an important role in areas of biology that investigate the relationships between biological entities such as species, viruses, and cells. A central task in the analysis of phylogenetic trees and networks is the quantification of the dissimilarity between them. Distances between phylogenetic trees that provide a measure of dissimilarity can be broadly classified into distances that are based on tree vector representations and those based on tree rearrangement operations [23]. While many of the former distances can be computed in polynomial time, the latter distances are typically NP-hard to compute. On the positive side, tree distances that are based on rearrangement operations provide a framework to explore and traverse spaces of phylogenetic trees since any tree in a given space can be transformed into any other tree of the same space by a finite number of operations such as nearest neighbour interchange (NNI), subtree prune and regraft (SPR), rooted subtree prune and regraft (rSPR), and tree bisection and reconnection (TBR) (see [24] for a summary and references therein for more details). In particular, the length of a shortest path between two phylogenetic trees in a given space of trees equals the distance between the two trees under the rearrangement operation that underlies the space.
@@
Computing distances that are based on tree rearrangement operations and related dissimilarity measures such as the minimum hybridisation number for two phylogenetic trees [4] remains an active area of research (e.g. [18, 22, 25]) despite the NP-hardness of the associated optimisation problems. Indeed, recent algorithmic progress facilitates computations that exactly calculate the aforementioned measures for data sets of remarkable size [26, 27]. Notably, the concept of agreement forests, which was first introduced by Hein et al. [13], underpins almost all mathematical and algorithmic advances in this area of research. Intuitively, an agreement forest of two binary phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ is a decomposition of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ into smaller and non-overlapping subtrees that have the same topology in ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Since the introduction of agreement forests almost 30 years ago, different variants of agreement forests have been used to characterise the rSPR distance between two rooted binary phylogenetic trees, the TBR distance between two unrooted binary phylogenetic trees, and the minimum hybridisation number of two rooted binary phylogenetic trees, as well as to establish related NP-hardness results [1, 4, 5]. Subsequent work has focussed on generalising agreement forests to collections of phylogenetic trees of arbitrary size that are not necessarily binary and on the development of fixed-parameter tractable and approximation algorithms (e.g. [8, 17, 21, 22, 28]). Additional developments in the context of agreement forests include a generalisation of agreement forests to relaxed agreement forests [2] and the exploitation of agreement forests to establish extremal results on the SPR, rSPR, and TBR distances [3, 9]. Part of the success of agreement forests is due to the fact that they replace the computation of a measure of dissimilarity between two phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ with the more static computation of an agreement forest such that the sought-after measure equates to the size of an optimal agreement forest for ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Moreover, agreement forests enable rigorous mathematical arguments that operate only on ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ without knowing the topology of any intermediate tree that lies on a shortest path between ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ in an associated space of trees.
@@
Since phylogenetic trees are somewhat limited in the type of biological processes that they can represent, rooted phylogenetic networks are increasingly being adopted to represent evolutionary relationships between biological entities whose past does not only include divergence events such as speciation but also convergence events such as lateral gene transfer or hybridisation. Inspired by tree rearrangement operations, several network rearrangement operations have recently been developed. For example, the subnet prune and regraft (SNPR) operation generalises rSPR to two rooted binary phylogenetic networks [6]. An SNPR operation either adds or deletes a reticulation edge (i.e., an edge that is directed into a vertex of in-degree two), or prunes and regrafts a subnetwork in the spirit of rSPR. Other network rearrangement operations for rooted phylogenetic networks have been developed and analysed in [11, 12, 14, 15, 16]. While a slightly modified framework of agreement forests can be used to compute the SNPR distance between a rooted binary phylogenetic tree and a rooted binary phylogenetic network [20], very little is known about how to compute distances that result from network rearrangement operations between two arbitrary rooted binary phylogenetic networks.
@@
$\frac{1}{2}m_{{\rm tc}}({\mathcal{N}},{\mathcal{N}}’)\leq d_{{\rm tc}}({\mathcal{N}},{\mathcal{N}}’)\leq m_{{\rm tc}}({\mathcal{N}},{\mathcal{N}}’).$ ’ ) .
@@
Our work relates to that of Klawitter [19], who has developed an alternative generalisation for two rooted binary phylogenetic networks. This generalisation for rooted networks gives rise to collections of agreement subgraphs that, in comparison with our generalisation, may have unlabelled leaves of in-degree one or two and that consequently do not resemble phylogenetic networks. In the same paper, Klawitter established bounds on the SNPR distance $d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ between two rooted binary phylogenetic networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ in terms of collections of agreement subgraphs whose number of unlabelled vertices of degree one is minimised. Without going into detail, this minimum number is referred to as $d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}^{\prime})$. In particular, Klawitter established the following theorem.
Theorem 1.2.
@@
[19, Corollary 5.5] Let ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ be two rooted binary phylogenetic networks on $X$. Then
$d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}’)\leq d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}’)\leq 6d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}^{\prime}).$
@@
While Theorem 1.2 applies to all rooted binary phylogenetic networks, it remains unknown whether or not the bounds are tight. Comparing Theorems 1.1 and 1.2, we note that $d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ in Theorem 1.2 refers to the minimum number of SNPR operations that are necessary to transform ${\mathcal{N}}$ into ${\mathcal{N}}^{\prime}$, while the quantity $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ in Theorem 1.1 does not only count SNPR operations but, additionally, weights them. More precisely, each SNPR operation that adds or deletes a reticulation edge is weighted one and each SNPR operation that prunes and regrafts a subnetwork is weighted two. Hence, $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ equates to the minimum sum of weights of SNPR operations that are needed to transform ${\mathcal{N}}$ into ${\mathcal{N}}^{\prime}$. Lastly, we note that although Klawitter’s generalisation and our definition of an agreement digraph differ, both definitions generalise agreement forests in the sense that, when applied to two rooted binary phylogenetic trees, they can be used to exactly compute their rSPR distance ([19, Proposition 4.2] and Proposition 6.6 of the present paper).
@@
The paper is organised as follows. In Section 2, we present basic notation and terminology for rooted phylogenetic networks. This is followed by an introduction of the new concepts of phylogenetic digraphs, which is the main definition building up to agreement digraphs, and their extensions in Section 3. Section 4 establishes several basic properties of extensions. Subsequently, in Section 5, we introduce the tree-child SNPR distance between two tree-child networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ and a maximum agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}$. Finally, in Section 6, we establish Theorem 1.1 and show that our bounds are tight.
2. Preliminaries
@@
This section provides notation and terminology that is used in the remainder of this paper. Throughout the paper, $X$ denotes a non-empty finite set. It is also worth noting that, except for rooted phylogenetic networks, general digraphs considered in this paper are not necessarily connected.
@@
Digraphs. Let $D$ be a digraph and let $v$ be a vertex in ${\mathcal{D}}$. We denote the edge set of $D$ by $E_{D}$ and say that $v$ is a tree vertex if $v$ has in-degree one and out-degree one or two, and that $v$ is a reticulation if $v$ has in-degree two and out-degree one. Furthermore, an edge $(u,v)$ in $D$ is a reticulation edge if $v$ is a reticulation and, otherwise, $(u,v)$ is a tree edge. Lastly, for two vertices $u$ and $v$ in $D$, we say that $u$ is a parent of $v$ and $v$ is a child of $u$ precisely if there is an edge $(u,v)$ in $D$.
Phylogenetic networks. A rooted binary phylogenetic network on $X$ is a rooted acyclic directed graph that satisfies the following properties:
@@
In keeping with the literature on distances between two phylogenetic networks, we allow parallel edges in rooted binary phylogenetic networks. Since all phylogenetic networks in this paper are rooted and binary, we simply refer to a rooted binary phylogenetic network on $X$ as a phylogenetic network on $X$. Now, let ${\mathcal{N}}$ be a phylogenetic network on $X$. The vertices of out-degree zero, that is the elements in $X$, are called leaves and $X$ is referred to as the leaf set of ${\mathcal{N}}$. If a phylogenetic network ${\mathcal{N}}$ has no reticulations, we call ${\mathcal{N}}$ a phylogenetic $X$-tree. Moreover, if ${\mathcal{N}}$ is a phylogenetic $X$-tree and $X$ contains exactly three elements, say $X=\{a,b,c\}$, then we refer to ${\mathcal{N}}$ as a triple and denote it by $(a,b,c)$ or, equivalently, $(b,a,c)$, if the underlying path joining the root of ${\mathcal{N}}$ and $c$ is vertex-disjoint from that joining $a$ and $b$.
Now, let ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ be two phylogenetic networks on $X$ with vertex and edge sets $V$ and $E$, and $V^{\prime}$ and $E^{\prime}$, respectively. We say that ${\mathcal{N}}$ is isomorphic to ${\mathcal{N}}^{\prime}$ if there is a bijection $\varphi:V\rightarrow V^{\prime}$ such that $\varphi(x)=x$ for all $x\in X$, and $(u,v)\in E$ if and only if $(\varphi(u),\varphi(v))\in E^{\prime}$ for all $u,v\in V$. If ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ are isomorphic, then we denote this by ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$.
Tree-child networks. Let ${\mathcal{N}}$ be a phylogenetic network on $X$. We say that ${\mathcal{N}}$ is tree-child if each non-leaf vertex has a child that is a tree vertex or a leaf. Moreover, we say that ${\mathcal{N}}$ contains a stack if there exist two reticulations that are joined by an edge and that ${\mathcal{N}}$ contains a pair of sibling reticulations if there exist two reticulations that have a common parent. The following well-known equivalence now follows from the definition of a tree-child network and will be freely used throughout the remainder of the paper.
Lemma 2.1.
Let ${\mathcal{N}}$ be a phylogenetic network. Then ${\mathcal{N}}$ is tree-child if and only if it has no stack, no sibling reticulations, and no pair of parallel edges.
The next lemma was established in [10, Lemma 7] and shows that the deletion of a reticulation edge of a tree-child network results in another tree-child network.
Lemma 2.2.
@@
3. Phylogenetic digraphs and their extensions

### v2

@@ Abstract
1 Introduction
@@
Phylogenetic trees and networks play an important role in areas of biology that investigate the relationships between biological entities such as species, viruses, and cells. A central task in the analysis of phylogenetic trees and networks is the quantification of the dissimilarity between them. Distances between phylogenetic trees that provide a measure of dissimilarity can be broadly classified into distances that are based on tree vector representations and those based on tree rearrangement operations [39]. While many of the former distances can be computed in polynomial time, the latter distances are typically NP-hard to compute. On the positive side and as summarised in Semple and Steel [40], tree distances that are based on rearrangement operations provide a framework to explore and traverse spaces of phylogenetic trees since any tree in a given space can be transformed into any other tree of the same space by a finite number of operations such as nearest neighbour interchange (NNI), subtree prune and regraft (SPR), rooted subtree prune and regraft (rSPR), and tree bisection and reconnection (TBR). In particular, the length of a shortest path between two phylogenetic trees in a given space of trees equals the distance between the two trees under the rearrangement operation that underlies the space.
@@
Computing distances that are based on tree rearrangement operations and related dissimilarity measures such as the minimum hybridisation number for two phylogenetic trees [5] remains an active area of research (e.g. [26, 38, 41]) despite the NP-hardness of the associated optimisation problems. Indeed, recent algorithmic progress facilitates computations that exactly calculate the aforementioned measures for data sets of remarkable size [43, 44]. Notably, the concept of agreement forests, which was first introduced by Hein et al. [16], underpins almost all mathematical and algorithmic advances in this area of research. Intuitively, an agreement forest of two binary phylogenetic trees $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ is a decomposition of $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ into smaller and non-overlapping subtrees that have the same topology in $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$. Since the introduction of agreement forests almost 30 years ago, different variants of agreement forests have been used to characterise the rSPR distance between two rooted binary phylogenetic trees, the TBR distance between two unrooted binary phylogenetic trees, and the minimum hybridisation number of two rooted binary phylogenetic trees, as well as to establish related NP-hardness results [1, 5, 6]. Subsequent work has focussed on generalising agreement forests to collections of phylogenetic trees of arbitrary size that are not necessarily binary and on the development of fixed-parameter tractable and approximation algorithms [9, 25, 33, 38, 45]. Additional developments in the context of agreement forests include a generalisation of agreement forests to relaxed agreement forests [3] and the exploitation of agreement forests to establish extremal results on the SPR, rSPR, and TBR distances [4, 11]. Part of the success of agreement forests is due to the fact that they replace the computation of a measure of dissimilarity between two phylogenetic trees $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ with the more static computation of an agreement forest such that the sought-after measure equates to the size of an optimal agreement forest for $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$. Moreover, agreement forests enable rigorous mathematical arguments that operate only on $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ without knowing the topology of any intermediate tree that lies on a shortest path between $\mathcal{T}$} and $\mathcal{T}$$^{\prime}$ in an associated space of trees.
@@
Since phylogenetic trees are somewhat limited in the type of biological processes that they can represent, rooted phylogenetic networks are increasingly being adopted to represent evolutionary relationships between biological entities whose past does not only include divergence events such as speciation but also convergence events such as lateral gene transfer or hybridisation. Inspired by tree rearrangement operations, several network rearrangement operations have recently been developed. For example, the subnet prune and regraft (SNPR) operation generalises rSPR to two rooted binary phylogenetic networks [7]. An SNPR operation either adds or deletes a reticulation edge (i.e., an edge that is directed into a vertex of in-degree two), or prunes and regrafts a subnetwork in the spirit of rSPR. Since the introduction of SNPR, the operation has been implemented in the Python package PhyloX [20], used in studies that, for example, reconstruct reassortment networks and involve a search through the space of rooted phylogenetic networks [35, 36, 37], and analysed mathematically with regards to the neighbourhood size [27] and properties of shortest length SNPR paths between two rooted phylogenetic networks [31]. Other network rearrangement operations for rooted phylogenetic networks [13, 15, 19, 21], unrooted phylogenetic networks [14, 17, 18, 22], and semi-directed networks [2, 34] have also been developed and analysed. Like SNPR, all such operations generalise a rearrangement operation for phylogenetic trees to phylogenetic networks. For excellent summaries of rearrangement operations for phylogenetic networks, we refer the reader to two PhD theses [23, 30]. Although a slightly modified framework of agreement forests can be used to compute the SNPR distance between a rooted binary phylogenetic tree and a rooted binary phylogenetic network [31], the question of how to compute the distances that result from network rearrangement operations between two arbitrary binary phylogenetic networks remains open.
@@
[…]
@@
The problem of finding an extension of an agreement digraph that optimally captures the commonalities between two rooted binary phylogenetic networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ may, at first sight, appear to be related to the problem of finding a maximum agreement subnetwork that $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ have in common [10, 24, 42]. However, upon further inspection, it becomes clear that the two problems are fundamentally different since a maximum agreement subnetwork for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ is not necessarily a component of an optimal agreement digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. On the other hand, our work is related to that of Klawitter [28, 29], who has developed an alternative generalisation of agreement forests for two rooted binary phylogenetic networks and for two unrooted binary phylogenetic networks. His generalisation for rooted networks gives rise to collections of agreement subgraphs that, in comparison with our generalisation, may have unlabelled leaves of in-degree one or two and that consequently do not resemble phylogenetic networks. In the same paper, Klawitter established bounds on the SNPR distance $d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ between two rooted binary phylogenetic networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ in terms of collections of agreement subgraphs whose number of unlabelled vertices of degree one is minimised. Without going into detail, this minimum number is referred to as $d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}^{\prime})$. In particular, Klawitter established the following theorem.
Theorem 2.
@@
[28, Corollary 5.5] Let $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ be two rooted binary phylogenetic networks on $X$. Then
$d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}’)\leqslant d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}’)\leqslant 6d_{\mathrm{AD}}({\mathcal{N}},{\mathcal{N}}^{\prime}).$
@@
While Theorem 2 applies to all rooted binary phylogenetic networks, it remains unknown whether or not the bounds are tight. Comparing Theorems 1 and 2, we note that $d^{*}_{\mathrm{SNPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ in Theorem 2 refers to the minimum number of SNPR operations that are necessary to transform $\mathcal{N}$} into $\mathcal{N}$$^{\prime}$, while the quantity $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ in Theorem 1 does not only count SNPR operations but, additionally, weights them. More precisely, each SNPR operation that adds or deletes a reticulation edge is weighted one and each SNPR operation that prunes and regrafts a subnetwork is weighted two. Hence, $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ equates to the minimum sum of weights of SNPR operations that are needed to transform $\mathcal{N}$} into $\mathcal{N}$$^{\prime}$. We discuss this further in the last section of the paper. Lastly, we note that although Klawitter’s generalisation and our definition of an agreement digraph differ, both definitions generalise agreement forests in the sense that, when applied to two rooted binary phylogenetic trees, they can be used to exactly compute their rSPR distance ([28, Proposition 4.2] and Proposition 15 of the present paper).
@@
The paper is organised as follows. In Section 2, we present basic notation and terminology for rooted phylogenetic networks. This is followed by an introduction of the new concepts of phylogenetic digraphs, which is the main definition building up to agreement digraphs, and their extensions in Section 3. Section 4 establishes several basic properties of extensions. Subsequently, in Section 5, we introduce the tree-child SNPR distance between two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ and a maximum agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$}. In Section 6, we establish Theorem 1 and show that our bounds are tight before we finish with some concluding remarks in Section 7.
2 Preliminaries
@@
This section provides notation and terminology that is used in the remainder of this paper. Throughout the paper, $X$ denotes a non-empty finite set. It is also worth noting that, except for rooted phylogenetic networks, the graphs that we consider in this paper are not necessarily connected. We start by introducing a broad class of directed acyclic graphs and several definitions that apply to this class. Subsequent definitions in this and the next section consider subclasses of these directed acyclic graphs.
@@
Directed acyclic graphs. Let $D$ be a directed acyclic graph. We allow parallel edges in $D$ and note that $D$ may have several vertices with in-degree zero. Furthermore, the undirected graph that underlies $D$ may contain more than one connected component. Let $V_{D}$ denote the vertex set of $D$, and let $E_{D}$ denote the edge set of $D$. We say that a vertex $v$ in $V_{D}$ is a tree vertex if $v$ has in-degree one and out-degree one or two, and that $v$ is a reticulation if $v$ has in-degree two and out-degree one. Furthermore, an edge $(u,v)$ in $D$ is a reticulation edge if $v$ is a reticulation and, otherwise, $(u,v)$ is a tree edge. Lastly, for two vertices $u$ and $v$ in $D$, we say that $u$ is a parent of $v$ and $v$ is a child of $u$ precisely if there is an edge $(u,v)$ in $D$.
Phylogenetic networks. Rooted phylogenetic networks generalise rooted phylogenetic trees to digraphs with underlying (but no directed) cycles. They allow vertices with in-degree greater than one, which represent non-treelike events such as hybridisation, lateral gene transfer, or recombination. Formally, a rooted binary phylogenetic network on $X$ is a connected directed acyclic graph with a single vertex of in-degree zero that satisfies the following properties:
@@
The vertices of $\mathcal{N}$} of out-degree zero are called leaves, and so $X$ is referred to as the leaf set ${\mathcal{L}}({\mathcal{N}})$ of $\mathcal{N}$}. In keeping with the literature on distances between two phylogenetic networks, we allow parallel edges in rooted binary phylogenetic networks. Since all phylogenetic networks in this paper are rooted and binary, we simply refer to a rooted binary phylogenetic network on $X$ as a phylogenetic network on $X$. Now, let $\mathcal{N}$} be a phylogenetic network on $X$. The vertices of out-degree zero, that is the elements in $X$, are called leaves and $X$ is referred to as the leaf set of $\mathcal{N}$}. If a phylogenetic network $\mathcal{N}$} has no reticulations, we call $\mathcal{N}$} a phylogenetic $X$-tree. Moreover, if $\mathcal{N}$} is a phylogenetic $X$-tree and $X$ contains exactly three elements, say $X=\{a,b,c\}$, then we refer to $\mathcal{N}$} as a triple if the underlying path joining the root of $\mathcal{N}$} and $c$ is vertex-disjoint from that joining $a$ and $b$ in which case, $(a,b,c)$ or, equivalently, $(b,a,c)$ denotes this triple.
Now, let $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ be two phylogenetic networks on $X$ with vertex and edge sets $V$ and $E$, and $V^{\prime}$ and $E^{\prime}$, respectively. We say that $\mathcal{N}$} is isomorphic to $\mathcal{N}$$^{\prime}$ if there is a bijection $\varphi:V\rightarrow V^{\prime}$ such that $\varphi(x)=x$ for all $x\in X$, and $(u,v)\in E$ if and only if $(\varphi(u),\varphi(v))\in E^{\prime}$ for all $u,v\in V$. If $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ are isomorphic, then we write ${\mathcal{N}}\cong{\mathcal{N}}^{\prime}$.
Tree-child networks. Let $\mathcal{N}$} be a phylogenetic network on $X$. We say that $\mathcal{N}$} is tree-child if each non-leaf vertex has a child that is a tree vertex or a leaf. Moreover, we say that $\mathcal{N}$} contains a stack if there exist two reticulations that are joined by an edge and that $\mathcal{N}$} contains a pair of sibling reticulations if there exist two reticulations that have a common parent. A phylogenetic network that is not tree-child is shown in Figure 1. In studying the mathematics that underlies phylogenetic networks, tree-child networks have been proven to be particularly successful because of their combinatorial properties that are often exploited to gain traction in establishing mathematical results. At the same time, these properties are not overly restrictive from a structural perspective in comparison to level-$1$ networks for example whose underlying cycles are pairwise vertex disjoint. For an overview of classes of phylogenetic networks, we refer the reader to Kong et al. [32].
[…]
Figure 1: A phylogenetic network on $X=\{1,2,3,4\}$ that is not tree-child because $u$ and $v$ form a stack, and $v$ and $v^{\prime}$ are a pair of sibling reticulations.
The following well-known equivalence follows from the definition of a tree-child network and will be freely used throughout the remainder of the paper.
Lemma 3.
Let $\mathcal{N}$} be a phylogenetic network. Then $\mathcal{N}$} is tree-child if and only if it has no stack, no pair of sibling reticulations, and no pair of parallel edges.
The next lemma was established by Döcker et al. [12, Lemma 7] and shows that the deletion of a reticulation edge of a tree-child network results in another tree-child network.
Lemma 4.
@@
3 Phylogenetic digraphs and their extensions